# Discrete equations from Bäcklund transformations of the fifth Painlevé equation

### v1

@@ Abstract
In this paper discrete equations are derived from Bäcklund transformations of the fifth Painlevé equation, including a new discrete equation which has ternary symmetry. There are two classes of rational solutions of the fifth Painlevé equation, one expressed in terms of the generalised Laguerre polynomials and the other in terms of the generalised Umemura polynomials, both of which can be expressed as Wronskians of Laguerre polynomials. Hierarchies of rational solutions of the discrete equations are derived in terms of the generalised Laguerre and generalised Umemura polynomials. It is known that there is nonuniqueness of some rational solutions of the fifth Painlevé equation. Pairs of nonunique rational solutions are used to derive distinct hierarchies of rational solutions which satisfy the same discrete equation.
@@ 1 Introduction
which we will refer to as PV.
@@
where $F$ is rational in ${\mathrm{d}}w/{\mathrm{d}}z$ and $w$ and analytic in $z$. The Painlevé transcendents, i.e. the solutions of the Painlevé equations, can be thought of as nonlinear analogues of the classical special functions. The Painlevé equations are regarded as being “completely integrable” and solvable using Riemann-Hilbert techniques [21]. The general solutions of the Painlevé equations are transcendental in the sense that they cannot be expressed in terms of known elementary functions and so require the introduction of a new transcendental function to describe their solution. However, it is well known that all the Painlevé equations, except PI, possess rational solutions, algebraic solutions and solutions expressed in terms of the classical special functions for special values of the parameters, see, e.g. [11, 33] and the references therein. These hierarchies are usually generated from “seed solutions” using the associated Bäcklund transformations and frequently can be expressed in the form of determinants.
@@
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$ [18, 25]. Hence letting $x_{n}=w(z;n+\lambda-\tfrac{1}{2})$, with $\alpha=n+\lambda-\tfrac{1}{2}$, then by eliminating $w^{\prime}(z;\alpha)$ we obtain another discrete equation, commonly known as alt-dPI
@@
which has PI (1.5) as its continuum limit [19]. The alt-dPI equation (1.9) appeared in 1981 in the work of Jimbo and Miwa [35, equation (5.3)], though it was not recognised to be a discrete Painlevé equation until the work of Fokas, Grammaticos and Ramani [19].
@@ 2 Bäcklund transformations of the fifth Painlevé equation
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, $\varepsilon_{j}=\pm 1$, $j=1,2,3$, independently, and for the parameters
@@
cf. [29, 30, 31], [33, §39] or [54, §32.7(v)].
Note that we are using the same notation as in [54, §32.7(v)] which slightly differs from that used in [29, 30, 31] or [33, §39].
@@
    $\alpha=\tfrac{1}{8}(\theta_{0}-\theta_{1}+\theta_{\infty})^{2},\qquad\beta=-\tfrac{1}{8}(\theta_{0}-\theta_{1}-\theta_{\infty})^{2},\qquad\gamma=1-\theta_{0}-\theta_{1},$
    see also [57], as are the Schlesinger transformations, which relate two solutions of the associated monodromy problem, for PV (1.2) [22, 46]. Further the discrete symmetries of PV (1.2), corresponding to the extended affine Weyl group of type $A_{3}^{(1)}$, act naturally on the root variables, see, for example, [13].
@@ 3 Deriving discrete Painlevé equations
To derive a discrete Painlevé equation from Bäcklund transformations of PV (1.2), we use a Bäcklund transformation $\mathcal{R}_{+}$, which relates a solution $w(z;a,b,c)$ to another solution $w_{+}(z;a_{+},b_{+},c_{+})$ and its inverse transformation $\mathcal{R}_{-}=\mathcal{R}_{+}^{-1}$, which relates $w(z;a,b,c)$ to a third solution $w_{-}(z;a_{-},b_{-},c_{-})$. Eliminating the derivative between the two Bäcklund transformations gives an algebraic relation between $w(z;a,b,c)$, $w_{+}(z;a_{+},b_{+},c_{+})$ and $w_{-}(z;a_{-},b_{-},c_{-})$, which is the discrete equation. Then setting $w(z;a,b,c)=w_{n}$, ${w_{\pm}}(z;a_{\pm},b_{\pm},c_{\pm})=w_{n\pm 1}$, with $(a,b,c)=(a_{n},b_{n},c_{n})$ and $(a_{\pm},b_{\pm},c_{\pm})=(a_{n\pm 1},b_{n\pm 1},c_{n\pm 1})$, gives the chain
@@
    It is essential that $\mathcal{R}_{-}$ is the inverse transformation of $\mathcal{R}_{+}$. It is possible to create an algebraic relation between three solutions of PV (1.2) using two Bäcklund transformations $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ when $\mathcal{R}_{2}\not=\mathcal{R}_{1}^{-1}$, but this won’t necessarily result in a discrete Painlevé equation. For example, in [33, §39], see also [30, 31], there are algebraic relations of three (and four) solutions of PV, but these are not discrete Painlevé equations.
@@
    cf. [54, §10]. Eliminating $\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}(z)$ between (3.2a) and (3.2b) yields the recurrence relation (3.2c). Letting $n\to n+1$ in (3.2a) and eliminating $J_{n+1}(z)$ using (3.2b) gives the differential equation (3.1). Thus the Bessel function $J_{n}(z)$ satisfies both the differential equation (3.1) and the difference equation (3.2c). An analogous situation occurs for the Painlevé equations, except PI (1.5).
@@
$\mathcal{R}_{4}=\mathcal{R}_{3}^{2}=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}^{2}\circ\mathcal{T}_{1,-1,1}.$
3.1 First discrete equation: asymmetric discrete Painlevé II
@@
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
$x_{n+1}+x_{n}$ $=\frac{4}{z}\frac{(2n+\lambda+1)y_{n}+\rho-\varphi}{1-y_{n}^{2}},$ (3.9a)
$y_{n}+y_{n-1}$ $=\frac{4}{z}\frac{(2n+\lambda)x_{n}+\rho+\varphi}{1-x_{n}^{2}},$ (3.9b)
@@ 3.2 Second discrete equation
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
$p_{n}+p_{n-1}$ $=\frac{4}{z}\,\frac{[\xi_{n}+(-1)^{n}\varphi]q_{n}+\rho}{1-q_{n}^{2}},$ (3.19a)
$q_{n+1}+q_{n}$ $=\frac{2}{z}\,\frac{\xi_{n+1}+\xi_{n}}{1-p_{n}},$ (3.19b)
@@
Remark 3.2.
Since the Bäcklund transformation $\mathcal{R}_{2}=\mathcal{T}_{-1,-1,-1}\circ\mathcal{T}_{-1,1,-1}$ involves a sequence of two Bäcklund transformations then it might be expected that the resulting transformation is quadratic in $\mathrm{d}{w}/\mathrm{d}{z}$. Generally if one applies the Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ followed by the Bäcklund transformation $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}$ then the resulting transformation is quadratic in $\mathrm{d}{w}/\mathrm{d}{z}$, but in some cases there is a simplification which happens in this case.
@@ 3.3 Third discrete equation: ternary discrete PI
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
    is a discrete PI [61, equation (3.6)], whilst the full equation, which was derived from a discrete dressing transformation [27], is a discrete form of PIV (1.10).
@@
    with $z=x+\mathrm{i}y\in\mathbb{C}$, and has the same unique solution as in [13]. Teodorescu et al. [68], see also [67, equation (26)], show that equation (3.32) arises in the theory of random normal matrices [74], which motivated the study in [17].
@@ 3.4 Fourth discrete equation: new equation with ternary symmetry
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}z$ and $a_{n}$, $b_{n}$ and $c_{n}$ are given by (3.25). Letting $w_{n}=1+1/x_{n}$ and $w_{n\pm 2}=1+1/x_{n\pm 2}$ gives
@@ 3.5 Differential-difference equations
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}=\tfrac{1}{2}(1-q_{n}^{2})\left\{\frac{2n+2\lambda+1}{q_{n+1}+q_{n}}-\frac{2n+2\lambda-1}{q_{n}+q_{n-1}}\right\}.$
@@
    $z\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}=-x_{n}(x_{n}+1)\left\{\frac{a_{n+1}}{x_{n+2}+x_{n}+1}-\frac{a_{n-1}}{x_{n}+x_{n-2}+1}-\right\}-b_{n}x_{n},$
@@
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}$ $=(1-q_{n}^{2})\left\{\frac{a_{2n+1}}{q_{n+1}+q_{n}}-\frac{a_{2n-1}}{q_{n}+q_{n-1}}\right\}-b_{2n}(q_{n}-1),$
    $z\frac{\mathrm{d}{p_{n}}}{{\mathrm{d}{z}}}$ $=(1-p_{n}^{2})\left\{\frac{a_{2n+2}}{p_{n+1}+p_{n}}-\frac{a_{2n}}{p_{n}+p_{n-1}}\right\}-b_{2n+1}(p_{n}-1),$
@@ 4 Rational solutions of PV
See Kitaev, Law and McLeod [40, Theorem 1.1]; also [33, Theorem 40.3].∎
@@
    Kitaev, Law and McLeod [40, Theorem 1.1] give four cases, though their cases (I) and (II) are related by the symmetry (4.2). Further Kitaev, Law and McLeod [40, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions, as illustrated in [12]; see also [3, 4].

### v2

@@ Abstract
In this paper discrete equations are derived from Bäcklund transformations of the fifth Painlevé equation, including a new discrete equation which has ternary symmetry. There are two classes of rational solutions of the fifth Painlevé equation, one expressed in terms of the generalised Laguerre polynomials and the other in terms of the generalised Umemura polynomials, both of which can be expressed as Wronskians of Laguerre polynomials. Hierarchies of rational solutions of the discrete equations are derived in terms of the generalised Laguerre and generalised Umemura polynomials. It is known that there is non-uniqueness of some rational solutions of the fifth Painlevé equation. Pairs of non-unique rational solutions are used to derive distinct hierarchies of rational solutions which satisfy the same discrete equation.
@@ 1 Introduction
which we will refer to as PV. In the case when $\delta=0$ in (1.1)
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=\left(\frac{1}{2w}+\frac{1}{w-1}\right)\left(\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right)^{\!\!2}-\frac{1}{z}\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}+\frac{(w-1)^{2}(\alpha w^{2}+{\beta})}{z^{2}w}+\frac{\gamma w}{z},$ (1.3)
which is known as the degenerate fifth Painlevé equation (deg-PV).
@@
where $F$ is rational in $\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}$ and $w$ and analytic in $z$. The Painlevé transcendents, i.e. the solutions of the Painlevé equations, can be thought of as nonlinear analogues of the classical special functions. The Painlevé equations are regarded as being “completely integrable” and solvable using Riemann-Hilbert techniques [25]. The general solutions of the Painlevé equations are transcendental in the sense that they cannot be expressed in terms of known elementary functions and so require the introduction of a new transcendental function to describe their solution. However, it is well known that all the Painlevé equations, except PI, possess rational solutions, algebraic solutions and solutions expressed in terms of the classical special functions for special values of the parameters, see, e.g. [14, 38] and the references therein. These hierarchies are usually generated from “seed solutions” using the associated Bäcklund transformations and frequently can be expressed in the form of determinants.
@@
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$ [22, 30]. Hence letting $x_{n}=w(z;n+\lambda-\tfrac{1}{2})$, with $\alpha=n+\lambda-\tfrac{1}{2}$, then by eliminating $w^{\prime}(z;\alpha)$ we obtain another discrete equation, commonly known as alt-dPI
@@
which has PI (1.6) as its continuum limit [23]. The alt-dPI equation (1.10) appeared in 1981 in the work of Jimbo and Miwa [41, equation (5.3)], though it was not recognised to be a discrete Painlevé equation until the work of Fokas, Grammaticos and Ramani [23].
@@ 2 Bäcklund transformations of the fifth Painlevé equation
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$, $\varepsilon_{j}=\pm 1$, $j=1,2,3$, independently, and for the parameters
@@
cf. [34, 35, 36], [38, §39] or [62, §32.7(v)].
Note that we are using the same notation as in [62, §32.7(v)] which slightly differs from that used in [34, 35, 36] or [38, §39].
@@
    $\alpha=\tfrac{1}{8}(\theta_{0}-\theta_{1}+\theta_{\infty})^{2},\qquad\beta=-\tfrac{1}{8}(\theta_{0}-\theta_{1}-\theta_{\infty})^{2},\qquad\gamma=1-\theta_{0}-\theta_{1},$ (2.5)
    see also [65], as are the Schlesinger transformations, which relate two solutions of the associated monodromy problem, for PV (1.2), see [26, 54]. Further the discrete symmetries of PV (1.2), corresponding to the extended affine Weyl group of type $A_{3}^{(1)}$, act naturally on the root variables, see, for example, [17]. Hence it is natural to define the Bäcklund transformations for PV (1.2) in terms of the parameters $a=\sqrt{2\alpha}$, $b=\sqrt{-2\beta}$ and $c=\gamma$, rather than $\alpha$, $\beta$ and $\gamma$.
@@ 3 Deriving discrete Painlevé equations
To derive a discrete Painlevé equation from Bäcklund transformations of PV (1.2), we use the Bäcklund transformation $\mathcal{R}_{+}$, which relates a solution $w(z;a,b,c)$ to another solution $w_{+}(z;a_{+},b_{+},c_{+})$ and its inverse transformation $\mathcal{R}_{-}=\mathcal{R}_{+}^{-1}$, which relates $w(z;a,b,c)$ to a third solution $w_{-}(z;a_{-},b_{-},c_{-})$. Eliminating the derivative between the two Bäcklund transformations gives an algebraic relation between $w(z;a,b,c)$, $w_{+}(z;a_{+},b_{+},c_{+})$ and $w_{-}(z;a_{-},b_{-},c_{-})$, which is the discrete equation. Then setting $w(z;a,b,c)=w_{n}$, ${w_{\pm}}(z;a_{\pm},b_{\pm},c_{\pm})=w_{n\pm 1}$, with $(a,b,c)=(a_{n},b_{n},c_{n})$ and $(a_{\pm},b_{\pm},c_{\pm})=(a_{n\pm 1},b_{n\pm 1},c_{n\pm 1})$, gives the chain
@@
    It is essential that $\mathcal{R}_{-}$ is the inverse transformation of $\mathcal{R}_{+}$. It is possible to create an algebraic relation between three solutions of PV (1.2) using two Bäcklund transformations $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ when $\mathcal{R}_{2}\not=\mathcal{R}_{1}^{-1}$, but this won’t necessarily result in a discrete Painlevé equation. For example, in [38, §39], see also [35, 36], there are algebraic relations of three (and four) solutions of PV, but these are not discrete Painlevé equations.
@@
    cf. [62, §10]. Eliminating $\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}(z)$ between (3.2a) and (3.2b) yields the recurrence relation (3.2c). Letting $n\to n+1$ in (3.2a) and eliminating $J_{n+1}(z)$ using (3.2b) gives the differential equation (3.1). Thus the Bessel function $J_{n}(z)$ satisfies both the differential equation (3.1) and the difference equation (3.2c). An analogous situation occurs for the Painlevé equations, except PI (1.6).
@@
$\mathcal{R}_{4}=\mathcal{R}_{3}^{2}=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}^{2}\circ\mathcal{T}_{1,-1,1},\qquad\mathcal{R}_{4}^{-1}=\mathcal{T}_{-1,1,1}^{2}.$
3.1 First discrete equation: asymmetric discrete PII
@@
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
$x_{n+1}+x_{n}$ $=\frac{4}{z}\frac{(2n+\lambda+1)y_{n}+\rho-\varphi}{1-y_{n}^{2}},\qquad y_{n}+y_{n-1}=\frac{4}{z}\frac{(2n+\lambda)x_{n}+\rho+\varphi}{1-x_{n}^{2}},$ (3.9)
@@
Remarks 3.2.
  1. 1.
    In the case when $\rho=\varphi=0$, equation (3.8) becomes
    $\frac{\mathrm{d}^{2}{q_{n}}}{{\mathrm{d}{z}}^{2}}=\frac{q_{n}}{q_{n}^{2}-1}\left\{\left(\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{(4n+\lambda)^{2}}{z^{2}}\right\}-\frac{1}{z}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}+\tfrac{1}{4}q_{n}(q_{n}^{2}-1),$ (3.12)
    which is equivalent to PV (1.2) with $\alpha=-\beta=\tfrac{1}{2}(n+\lambda)^{2}$ and $\gamma=0$. Making the transformation $q_{n}=1/\sqrt{1-u_{n}}$ with $z=2\sqrt{x}$ in equation (3.12) gives deg-PV (1.3) with $\alpha=0$, $\beta=-2(n+\lambda)^{2}$ and $\gamma=\tfrac{1}{2}$. Solutions of deg-PV (1.3) are related to those of PIII (3.10), cf. [38, Theorem 34.2]. Solutions of equation (3.12), in terms of Bessel functions, are discussed in [15, 39].
  2. 2.
    Equation (3.12) arises as a reduction of the two-dimensional complex sine-Gordon equation, also known as the Pohlmeyer-Lund-Regge model [48, 49, 69], in quantum systems [10, 11, 12], in relativity [29] and in coefficients of the three-term recurrence relation for orthogonal polynomials with respect to the weight $w(\theta)=\exp({t\cos\theta})$ on the unit circle, see [80, equation (3.13)]. The orthogonal polynomials for this weight are related to unitary random matrices [68].
@@ 3.2 Second discrete equation
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
$p_{n}+p_{n-1}$ $=\frac{4}{z}\,\frac{[\xi_{n}+(-1)^{n}\varphi]q_{n}+\rho}{1-q_{n}^{2}},\qquad q_{n+1}+q_{n}=\frac{2}{z}\,\frac{\xi_{n+1}+\xi_{n}}{1-p_{n}},$ (3.20)
@@
Remark 3.3.
Since the Bäcklund transformation $\mathcal{R}_{2}=\mathcal{T}_{-1,-1,-1}\circ\mathcal{T}_{-1,1,-1}$ involves a sequence of two Bäcklund transformations then it might be expected that the resulting transformation is quadratic in $\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}$. Generally if one applies the Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ followed by the Bäcklund transformation $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}$ then the resulting transformation is quadratic in $\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}$, but in some cases there is a simplification which happens in this case.
@@ 3.3 Third discrete equation: ternary discrete PI
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
@@
    is a discrete PI [70, equation (3.6)], whilst the full equation, which was derived from a discrete dressing transformation [32], is a discrete form of PIV (1.11).
@@
    with $z=x+\mathrm{i}y\in\mathbb{C}$, and has the same unique solution as in [17]. Teodorescu et al. [77], see also [76, equation (26)], show that equation (3.33) arises in the theory of random normal matrices [83], which motivated the study in [21].
@@ 3.4 Fourth discrete equation: new equation with ternary symmetry
with $^{\prime}\equiv\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$ and $a_{n}$, $b_{n}$ and $c_{n}$ are given by (3.26). Letting $w_{n}=1+1/x_{n}$ and $w_{n\pm 2}=1+1/x_{n\pm 2}$ gives
@@ 3.5 Differential-difference equations
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}=\tfrac{1}{2}(1-q_{n}^{2})\left(\frac{2n+2\lambda+1}{q_{n+1}+q_{n}}-\frac{2n+2\lambda-1}{q_{n}+q_{n-1}}\right).$
@@
    $z\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}=-x_{n}(x_{n}+1)\left(\frac{a_{n+1}}{x_{n+2}+x_{n}+1}-\frac{a_{n-1}}{x_{n}+x_{n-2}+1}\right)-b_{n}x_{n},$
@@
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}$ $=(1-q_{n}^{2})\left(\frac{a_{2n+1}}{q_{n+1}+q_{n}}-\frac{a_{2n-1}}{q_{n}+q_{n-1}}\right)-b_{2n}(q_{n}-1),$
    $z\frac{\mathrm{d}{p_{n}}}{{\mathrm{d}{z}}}$ $=(1-p_{n}^{2})\left(\frac{a_{2n+2}}{p_{n+1}+p_{n}}-\frac{a_{2n}}{p_{n}+p_{n-1}}\right)-b_{2n+1}(p_{n}-1),$
@@ 4 Rational solutions of PV
See Kitaev, Law and McLeod [46, Theorem 1.1]; also [38, Theorem 40.3].∎
@@
    Kitaev, Law and McLeod [46, Theorem 1.1] give four cases, though their cases (I) and (II) are related by the symmetry (4.2). Further Kitaev, Law and McLeod [46, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions, as illustrated in [16]; see also [3, 4].